20 gm of a MgCO3 samle decomposes on heatinh to give CO2 and 8.0fm MgO what vill the %age purity of MgCO3 in the sample

Grade:12th pass

1 Answers

Vikas TU
4 years ago
Hiii 
Mg Co3 => Mgo + Co2 
8 g of mg is formed 
let y gm of MgCo3 is remain 
y = 84*8/40 = 16.8 g 
% Purity of Mg co3 = 16.8 * 100 / 20 = 84%